INAF queue observing, 2020 Jan 27 - Jan 28

Observers: A. Rossi, S. Paiano
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LUCI-1 & LUCI-2

During the afternoon, around 24:00UT the weather is windy. The support astronomer said us that the telescopes need ventilation to stabilize the temperature of the mirrors.

At 0:45UT, we are opening the dome! The weather is a little bit cloudy and the humidity is increasing (~72%).

Summary

Night duration (12 degree - twilight) 1:30 - 07:30 UT = 06 h 00m = 21600 seconds

Observing time: 05h 00min (83.3%)

Weather time loss: 45min (12.5%)

technical time loss 15min (4.2%)

Data Summary

Calibrations

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations Darks LUCI1-2 5 NDIT x 3.0 DIT , LIR/INT     for different programs
Calibrations Darks LUCI1-2 6 NDIT x 5.0 DIT , LIR/INT     for different programs
Calibrations Darks LUCI1-2 1 NDIT x 200 DIT , MER/NORM     for different programs
Calibrations Darks LUCI1-2 1 NDIT x 300 DIT , MER/NORM     for different programs
Calibrations Darks LUCI1-2 2 NDIT x 30.0 DIT , LIR/INT Lamps OFF     for different programs
Calibrations Darks LUCI1-2 2 NDIT x 30.0 DIT , LIR/INT Lamps ON (spec flat)     for different programs
Calibrations Darks LUCI1-2 1 NDIT x 20.0 DIT , LIR/NORM Argon Arc     for different programs
Calibrations Darks LUCI1-2 1 NDIT x 60.0 DIT , LIR/NORM     for different programs
Calibrations Flat & Arcs LUCI1-2 LS-K210 2.24 0.25"     Teresa Giannini
Calibrations Flat & Arcs LUCI1-2 Mask-GargiuloVAN003     Adriana Gargiulo
Calibrations Flat & Arcs LUCI1-2 Mask-GargiuloVAN001     Adriana Gargiulo
Calibrations Flat & Arcs LUCI1-2 LS-zj200 1.17 1.00"     Andrea Rossi
Calibrations Bias MODS1-2 Bias8k     for different programs
Calibrations Bias MODS1-2 Bias3k     for different programs
Calibrations Flat & Arcs LUCI1-2 LS-hk200 1.93 1.00"     Andrea Rossi - taken during the morning
Calibrations Flats MODS1-2 Slit 1.2" - Dual grating     for different programs
Calibrations Arcs MODS1-2 Slit 1.2" - Dual grating     for different programs
Calibrations Flats MODS1-2 Slitless     for different programs

Scientific Targets

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Telluric HIP10185 LUCI1-2 zj200   luci1.20200128.0078-0080 / luci2.20200128.0068-0071 for Gargiulo
Telluric HIP20379 LUCI1-2 zj200 1.00" - luci1.20200128.00 / luci2.20200128.00 for Andrea Rossi
Rossi SN2018ibb LUCI2 zj200 1.00" 3600 sec luci2.20200128 0087-0104 for Andrea Rossi - LUCI1 is unavailable completed
Rossi SN2018ibb LUCI2 hk200 1.00" 3600 sec luci2.20200128.0105-0118 for Andrea Rossi - LUCI1 is unavailable completed
Telluric HD52266 LUCI2 LS-K210 2.24 0.25" - luci2.20200128.0123-0126 for Giannini - LUCI1 is unavailable
Giannini V960-mon LUCI2 LS-K210 2.24 0.25" 1800 sec luci2.20200128.0131- for Giannini - LUCI1 is unavailable completed
Telluric HD52266 LUCI2 LS-K210 2.24 0.25" - luci2.20200128.0165-0168 for Giannini - LUCI1 is unavailable
Giannini V900-mon LUCI2 LS-K210 2.24 0.25" 0 sec luci2.20200128.0131- for Giannini - LUCI1 is unavailable - Lost for bad weather

Details

01:34 We are pointing the Telluric for Gargiulo's program. MOS data + zj200 1.00"
  • acq = luci1.20200128.0074-76= =luci2.20200128.0065-67 = seeing=>1.5'
  • 01:52 seeing 1.6"
  • 01:52 sci = luci1.20200128.0078-0080 luci2.20200128.0068-0071 =

02:00 We are pointing the Telluric for DDT Andrea's program. LS data with zj200 and HK200 1.00"
  • acq = luci1.20200128.0081-83= =luci2.20200128.0072-75 = seeing=1.5
  • Seeing=1.3-1.5"
  • 02:10 sci = luci1.20200128.0084-0091 luci2.20200128.0076-0082 =

02:16 We are pointing the SN2018ibb object of DDT Andrea's program . LS data with zj200 1.00"
  • acq = luci1.20200128.0092-95- = =luci2.20200128.0083-86- = seeing=1.4
  • Seeing=1.4"
  • 02:XX sci = luci1.20200128.0096-0097 luci2.20200128.0087-on-going =

The first and second spectrum of LUCI1 are bad, because it is out of focus. = 10 min Time loss from 02:40 to 02:50=

LUCI1 is unavailable from 02:50 due to a focus problem. We continue the observations using only LUCI2.

02:50 We are still pointing the SN2018ibb object of DDT Andrea's program but only using LUCI2 + zj200 1.00". LS data.
  • Seeing=1.5"
  • 02:50 sc = luci2.20200128.0087-0104 =
  • 03:05 seeing 1.3" on DIMM humidity is 80-90% since 1 hour and windy

03:55 We are still pointing the SN2018ibb object of DDT Andrea's program changing the setting. LS data with hk200-1.93 1.00"
  • acq = luci2.20200128.0083-86- = seeing=1.5
  • Seeing=1.5"-1.8" , Temp=-5 , Hum=85%, Wind=12.5m/s, no clouds
  • 03:55 sci = luci2.20200128.0105-0118 Probably the last two spectra should be rejected due to bad seeing (>2.0")=

Seeing around 1.8" from 04:40UT. Seeing>2.0"-2.5" from 04:45UT and it is increasing.

At 04:50 we paused the observation to disconnet LUCI1. We resume it after 5 minutes

05:00 We changed the source and we are observing the Telluric HD52266 for Giannini's program.
  • acq = luci2.20200128.0119-0122- = seeing=>2.0"
  • 05:10 seeing 1.9"
  • 05:10 sci = luci2.20200128.0123-0126 =

At 05:10UT, there are some strong wind gusts.

05:15 We are observing the first target of Giannini, the star V960-mon. LS data with LS-K210 2.24 0.25"
  • acq = luci2.20200128.0127-0130 = seeing=1.8
  • Seeing=1.8"-1.5" , Temp=-6 , Hum=87%, Wind=16.5m/s, no clouds
  • 05:28 sci = luci2.20200128.0131-0160 =

During the observation the seeing is unstable, spanning from 1.2 to 1.5. The humidity is increasing up to 90%.

06:13 We observe again the Telluric HD52266 for Giannini's program.
  • acq = luci2.20200128.0161-0164 = seeing=1.1-1.5"
  • 06:22 seeing variable and around 1.2", Temp=-6 , Hum=92%, Wind=14.9m/s, no clouds
  • 06:25 sci = luci2.20200128.0165-0168 =

06:30 We are observing the second target of Giannini, the star V900-mon. LS data with LS-K210 2.24 0.25"
  • acq = luci2.20200128.0169-0171 = seeing=1.2"
  • Seeing=1.2" , Temp=-6 , Hum=94%, Wind=16.5m/s, no clouds
  • 06:37 sci = no data =

At 06:37UT, the humidity is 95.5% (T=-6C). We have to close the dome for high humidity.

Finally, we took some calibration:
  • Bias 8k for MODS1-2
  • Bias 3k for MODS1-2

At 14.14UT we start to take some calibrations for MODS1-2:
  • Flats for MODS1-2 for slit 1.2" - dual grating
  • Flats & arcs for LS-hk200 1.93 1.00" configuration for LUCI2
  • Arcs for MODS1-2 for slit 1.2" - dual grating
  • Flats for MODS1-2 for slitless

CALIBRATIONS took from Dave for us the day before
DARKS
  • Darks_giannini_V900V960telluric_bino.xml
  • LUCI1 LUCI2 Comments________________
  • 0001 0005 0001 0005 5x[5x12.0s] LIR/INT
  • 0006 0010 0006 0010 5x[6x5.0s] LIR/INT
  • 0011 0015 0011 0015 5x[2x30.0s] LIR/INT

  • darks_rossiDDTspec+ima_bino.xml
  • LUCI1 LUCI2 Comments________________
  • 0016 0020 0016 0020 5x[5x3.0s] LIR/INT
  • 0021 0025 0021 0025 5x[1x200.0s] MER/NORM

Topic revision: r8 - 07 Feb 2020, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback